# Adjoint motives of modular forms and the Tamagawa number conjecture

### v1

@@ 3.2. Action on curves
where the top row is obtained by applying $\varepsilon_{h}^{-1}$ to (3), the bottom row is (3) with $N^{\prime}$ instead of NN, and the vertical arrows are given by the canonical $\varepsilon_{h}^{-1}{\mathcal{O}}_{\bar{E}}$-linear maps in each degree. Applying ${\bf R}^{\bullet}\bar{s}^{\prime}_{*}$ and using the natural transformation $\chi_{h}^{-1}\bar{s}_{*}\to\bar{s}^{\prime}_{*}\varepsilon_{h}^{-1}$ gives a commutative diagram
@@
Fix positive integers $k\geq 2$, and NN, $N^{\prime}\geq 3$. Let $g$ be an element in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$ such that $g^{-1}U_{N^{\prime}}g\subseteq U_{N}$.
@@ 3.4. Relation with the action on modular forms
Further let $r\in{\mathbb{Z}}$. Then the action of rIrI on both spaces are multiplication by $r^{k-2}$. This proves the proposition. ∎
@@ 3.6. Compatibility with changing of levels
Fix $g\in M_{2}({\mathbb{Z}})\cap\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$. Let $N,N^{\prime},M$ and $M^{\prime}$ be integers such that $N|M,N^{\prime}|M^{\prime},g^{-1}U_{N^{\prime}}g\subset U_{N}$ and $g^{-1}U_{M^{\prime}}g\subset U_{M}$. Then we have $g]:{\mathcal{M}}_{N}\to{\mathcal{M}}_{N^{\prime}}$ and $g]:{\mathcal{M}}_{M}\to{\mathcal{M}}_{M^{\prime}}$. Let II be the identity matrix in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$. From $N|M$ we get $I^{-1}U_{M}I\subset U_{N}$ and we have
@@
4. The premotivic structure for forms of level NN and character $\psi$
@@ 4.1. $\sigma$-constructions
Because of the lemma, we will suppress NN in the notations by using ${\mathcal{M}}(\sigma)_{c}$, ${\mathcal{M}}(\sigma)_{!}$ and $M(\sigma)$ when there is no danger of confusion.
@@ 4.2. Action of double cosets
To see that this is possible, first choose NN divisible by 44 so that $S_{N}^{K}\subset S$ and the second condition holds, then choose $N^{\prime}$ so that the third condition holds, and then for each $\ell\not\in S$, replace $N^{\prime}$ by $N^{\prime}/\ell^{r}$ where $\ell^{r}||N^{\prime}$. The condition that $S_{N}^{K}\subset S$ and $g_{\ell}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{\ell})$ ensures that we still have $g^{-1}U_{N^{\prime}}g\subset U_{N}.$ Note that $N|N^{\prime}$ so $N^{\prime}\geq 3$, $U_{N^{\prime}}\subset U_{N}\subset\ker\sigma\cap\ker\sigma^{\prime}$ and $S_{N}^{K}\subset S_{N^{\prime}}^{K}\subset S$. Then from the last section, $g$ gives a map $g]:{\mathcal{M}}_{N}\to{\mathcal{M}}_{N^{\prime}}$. Further $U^{\prime}$ acts on ${\mathcal{M}}_{N^{\prime}}$. Let $U^{\prime}=\coprod_{i}g_{i}U_{1}^{\prime}$ be a coset decomposition. We define a morphism $U^{\prime}gU]_{\tau,N,N^{\prime}}$ by restricting the map
@@ 4.3. Compositions
We will give a formula for the composition $U^{\prime\prime}g^{\prime}U^{\prime}]_{\tau^{\prime},N^{\prime},N^{\prime\prime},\sharp}\circ[U^{\prime}gU]_{\tau,N,N^{\prime},\sharp}$ for $\sharp\in\{\phi,c,!\}$. To simplify the notation, we will suppress NN, $N^{\prime}$ and $\sharp$ when there is no danger of confusion.
@@
Using Lemma 3.9 and Proposition 4.4, we conclude that $U^{\prime}gU]$ is independent of the choice of NN and $N^{\prime}$. More precisely, if $U^{\prime}gU]_{\tau,M,M^{\prime}}$ is also defined, then $U^{\prime}gU]_{\tau,N,N^{\prime}}^{S}=[U^{\prime}gU]_{\tau,M,M^{\prime}}^{S}$ when $S\supset S_{\operatorname{{\mathop{\rm lcm}\nolimits}}(N^{\prime},M^{\prime})}$.
$4.$ Pairing on $\sigma$-constructions
@@
For NN, $N^{\prime}$ and $g$ as above, we have the commutative diagram
@@
Suppose NN is such that $N\geq 3$, $U_{N}\subset U$, the conductor of $\psi$ divides NN. Then the morphism ${\mathcal{M}}_{F}\to{\mathcal{H}}$ induces an isomorphism
@@
We now construct the pairings by defining the duality homomorphisms. Given a representation $\sigma:U\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{{\mathcal{O}}_{K}}{\mathcal{V}}$, let $\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits})$ denote the representation defined by the action of $U$ on ${\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{V}},{\mathcal{O}}_{\psi})$. For $N\geq 3$ such that $U_{N}\subset\ker\sigma$ and the conductor of $\psi$ divides NN, Lemma (4.7) and the duality homomorphism $\delta^{L}_{N}:{\mathcal{M}}_{N,c}\to{\mathop{\rm Hom}\nolimits}({\mathcal{M}}_{N},{\mathcal{T}}^{S_{N},\otimes(k-1)})$ from §2.5 give a morphism
@@
We will therefore omit the subscript NN in the notation for the morphisms $\bar{\delta}^{L}$. (Note however that $\delta^{L}_{N}$ depends on NN, and $\delta^{L}_{N,!}$ depends on NN only in the normalization.)
@@
To do this, we give an alternate description of the Betti realizations and pairings which is independent of NN. and define a perfect pairing on the new Betti realizations. Given a sufficiently small $U$ and a $\sigma:U\to\operatorname{{\mathop{\rm Aut}\nolimits}}_{{\mathcal{O}}_{K}}{\mathcal{V}}$, define a locally constant sheaf ${\mathcal{F}}(\sigma)$ on $G_{\mathbb{Q}}\backslash G_{\mathbb{A}}/UU_{\infty}$ by ${\mathcal{F}}(\sigma)=G_{\mathbb{Q}}\backslash(G_{\mathbb{A}}\times{\mathcal{V}})/UU_{\infty}$. Define
@@
4.5. Premotivic structure of level NN and character $\psi$
@@
for all ${\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in U_{0}(N)$. It is straightforward to check that this is precisely the space $M_{k}(N,\psi)$ of classical modular forms of weight $k$, level NN and character $\psi$, i.e., the set of forms $f\in M_{k}(\Gamma(M))$ such that
@@ 5.1. Hecke actions
Fix a rational prime pp. By Lemma 4.2, $U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}_{p}U]_{\psi(p_{p})^{-1}}$ defines an endomorphism of ${\mathcal{M}}(N,\psi)_{M,!}$ regarded as an object in ${\mathcal{P}M}^{S_{pN}}$. This induces an endomorphism
@@ 5.2. Compatibility of the Hecke action with the pairings
Since NN is invertible in $\mathcal{O}$}, $UwU]_{\omega}$ is an isomorphism in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$ and
@@ 5.3. The action of Hecke rings
Fix a finite prime pp. Recall that ${\mathcal{M}}(N,\psi)_{M,!}$ is in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$ and, by Lemma 4.2, $T_{p}:{\mathcal{M}}(N,\psi)_{M,!}\to{\mathcal{M}}(N,\psi)_{pM,!}$ is defined in ${\mathcal{P}M}^{S_{pM}^{K}}_{K}={\mathcal{P}M}^{S_{pN}^{K}}_{K}$. If $p|Nk!$, then $S_{pN}=S_{N}$. So $T_{p}$ is a morphism in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$.
@@ 6.3. $\Sigma$-variations for non-exceptional primes
In Proposition 4.4 take $U=U^{\Sigma_{0}},U^{\prime}=U^{\Sigma_{1}},U^{\prime\prime}=U^{\Sigma_{2}},g={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{m_{1}}}\end{array}\right)$}},g^{\prime}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{m_{2}}}\end{array}\right)$}}$ and $\tau=\tau^{\prime}={\mathop{\rm id}\nolimits}$. Then we have $U^{\prime}_{1}=U^{\prime}$. So in the double coset decomposition $U^{\prime}=\coprod_{k}U^{\prime}_{2}u_{k}U^{\prime}_{1}$ there is only one $u_{k}$ which we can take II. Further $W_{1}=W_{1}^{\prime}=U^{\prime\prime}$ and $\tau_{1}=1$. This gives us the desired equation.
@@
For this we first display the following elementary relations. Fix a prime number pp. For each integer $i\geq 0$, let $V_{i}=\{{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})|c\equiv 0\mod p^{i}\}.$ Let $A_{t}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p^{t}}}\end{array}\right)$}},t=0,1,2$. Let $B_{0}=I,B_{1}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}},B_{2}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}$. Then
@@
Thus to prove Proposition 6.3, we only need to prove it in the case when $\Sigma$ contains one prime pp. Let $\Sigma_{0}=\phi$ and $\Sigma_{1}=\Sigma$. We consider three cases depending on $\delta_{p}=0,1$ or 22.
@@
When $q=p$, by (41) and (42), the double coset product of the local factor at primes other than $p$ involves II and is obviously zero. The local factor at $p$ is
@@
We consider the action of $\gamma=\epsilon_{1}\circ\phi_{1}+\epsilon_{p}\circ\phi_{p}$ on $f$ as a classical modular form. First $\epsilon_{1}\circ\phi_{1}=[U{{}^{\Sigma}}IU]_{1}$ acts on ${\mathcal{M}}(N,\psi)_{M}\subseteq{\mathcal{M}}_{M}\otimes{\mathcal{V}}$ by $I\otimes 1$, where $M\geq 3$ is chosen such that $N|M$ and $S_{M}=S_{N}$. The action corresponding to II on the classical forms is II. Since $f$ has rational coefficients, its image in $\oplus_{t\in({\mathbb{Z}}/M{\mathbb{Z}})^{\times}}M(M,\psi)$ is of the form $(f_{0}(z))_{t}$ with $f_{0}$ independent of $t$. Then, in the notations of Lemma 3.2 and Proposition 3.7, for each $t^{\prime}\in({\mathbb{Z}}/pM{\mathbb{Z}})^{\times}$ we have $\kappa(t^{\prime})=t^{\prime}\mod M$ and $\tilde{\gamma}:E_{pM,t^{\prime}}\to E_{M,t}$ defined by $\gamma=\gamma_{t^{\prime}}=I.$ So $(f_{0}(z))_{t}$ is sent to $(f_{0}(z))_{t^{\prime}}$, $t^{\prime}\in({\mathbb{Z}}/pM{\mathbb{Z}})^{\times}$ by $\epsilon_{1}\circ\phi_{1}$. Next $U{{}^{\Sigma}}=U{{}^{\Sigma}}\cap({\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}^{-1}$. So $\epsilon_{p}=p^{1-k}[U{{}^{\Sigma}}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}U]_{1}$ acts on ${\mathcal{M}}_{pM}\otimes{\mathcal{V}}$ by ${\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}\otimes 1$. The corresponding action on $\oplus_{t}M_{k}(\Gamma_{0}(pN))$ is given by taking, for each $t^{\prime}\in({\mathbb{Z}}/p^{2}M{\mathbb{Z}})^{\times}$, $\kappa(t^{\prime})=t^{\prime}\mod pM$ and $\tilde{\gamma}_{t^{\prime}}:E_{p^{2}M,t^{\prime}}\to E_{pM,t}$ given by $\gamma_{t^{\prime}}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p^{-1}}}\end{array}\right)$}}.$ So $\epsilon_{p}$ sends $(f_{0})_{t}\in\oplus_{t}M_{k}(\Gamma_{0}(pM))$ to $(f_{0}(pz))_{t^{\prime}}$. Since $\phi_{p}=-T_{p}$, $\epsilon_{p}\circ\phi_{p}$ sends $(f_{0}(z))_{t}$ to $(-f_{0}{}_{|T_{p}})_{t^{\prime}}$. So by [Miy, Lemma 4.6.5], $\gamma$ sends $(f_{0}(z))_{t}$ to $(f_{0}{{}^{\Sigma}}(z))_{t^{\prime}}$, as is needed.
@@
When $q\neq p$, the relation $T_{q}\gamma=\gamma T_{q}$ is proved in the same way as in Case 2. When $q=p$, we only need to consider the local factor of $T_{p}\gamma$ at the prime pp. The double coset product of the local factor is
@@
As in the proof of Proposition 6.3, we can again specialize to the case when $\Sigma$ has a single prime pp. Then we only need to consider the place at pp.

### v2

@@ 3.3. Action on the premotivic structures
where the top row is obtained by applying $\varepsilon_{h}^{-1}$ to (3), the bottom row is (3) with $N^{\prime}$ instead of $N$, and the vertical arrows are given by the canonical $\varepsilon_{h}^{-1}{\mathcal{O}}_{\bar{E}}$-linear maps in each degree. Applying ${\bf R}^{\bullet}\bar{s}^{\prime}_{*}$ and using the natural transformation $\chi_{h}^{-1}\bar{s}_{*}\to\bar{s}^{\prime}_{*}\varepsilon_{h}^{-1}$ gives a commutative diagram
@@
Fix positive integers $k\geq 2$, and $N$, $N^{\prime}\geq 3$. Let $g$ be an element in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$ such that $g^{-1}U_{N^{\prime}}g\subseteq U_{N}$.
@@ 3.4. Relation with the action on modular forms
Further let $r\in{\mathbb{Z}}$. Then the action of $rI$ on both spaces are multiplication by $r^{k-2}$. This proves the proposition. ∎
@@ 3.6. Compatibility with changing of levels
Fix $g\in M_{2}({\mathbb{Z}})\cap\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$. Let $N,N^{\prime},M$ and $M^{\prime}$ be integers such that $N|M,N^{\prime}|M^{\prime},g^{-1}U_{N^{\prime}}g\subset U_{N}$ and $g^{-1}U_{M^{\prime}}g\subset U_{M}$. Then we have $g]:{\mathcal{M}}_{N}\to{\mathcal{M}}_{N^{\prime}}$ and $g]:{\mathcal{M}}_{M}\to{\mathcal{M}}_{M^{\prime}}$. Let $I$ be the identity matrix in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$. From $N|M$ we get $I^{-1}U_{M}I\subset U_{N}$ and we have
@@
4. The premotivic structure for forms of level $N$ and character $\psi$
@@ 4.1. $\sigma$-constructions
Because of the lemma, we will suppress $N$ in the notations by using ${\mathcal{M}}(\sigma)_{c}$, ${\mathcal{M}}(\sigma)_{!}$ and $M(\sigma)$ when there is no danger of confusion.
@@ 4.2. Action of double cosets
To see that this is possible, first choose $N$ divisible by $4$ so that $S_{N}^{K}\subset S$ and the second condition holds, then choose $N^{\prime}$ so that the third condition holds, and then for each $\ell\not\in S$, replace $N^{\prime}$ by $N^{\prime}/\ell^{r}$ where $\ell^{r}||N^{\prime}$. The condition that $S_{N}^{K}\subset S$ and $g_{\ell}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{\ell})$ ensures that we still have $g^{-1}U_{N^{\prime}}g\subset U_{N}.$ Note that $N|N^{\prime}$ so $N^{\prime}\geq 3$, $U_{N^{\prime}}\subset U_{N}\subset\ker\sigma\cap\ker\sigma^{\prime}$ and $S_{N}^{K}\subset S_{N^{\prime}}^{K}\subset S$. Then from the last section, $g$ gives a map $g]:{\mathcal{M}}_{N}\to{\mathcal{M}}_{N^{\prime}}$. Further $U^{\prime}$ acts on ${\mathcal{M}}_{N^{\prime}}$. Let $U^{\prime}=\coprod_{i}g_{i}U_{1}^{\prime}$ be a coset decomposition. We define a morphism $U^{\prime}gU]_{\tau,N,N^{\prime}}$ by restricting the map
@@ 4.3. Compositions
We will give a formula for the composition $U^{\prime\prime}g^{\prime}U^{\prime}]_{\tau^{\prime},N^{\prime},N^{\prime\prime},\sharp}\circ[U^{\prime}gU]_{\tau,N,N^{\prime},\sharp}$ for $\sharp\in\{\phi,c,!\}$. To simplify the notation, we will suppress $N$, $N^{\prime}$ and $\sharp$ when there is no danger of confusion.
@@
Using Lemma 3.9 and Proposition 4.4, we conclude that $U^{\prime}gU]$ is independent of the choice of $N$ and $N^{\prime}$. More precisely, if $U^{\prime}gU]_{\tau,M,M^{\prime}}$ is also defined, then $U^{\prime}gU]_{\tau,N,N^{\prime}}^{S}=[U^{\prime}gU]_{\tau,M,M^{\prime}}^{S}$ when $S\supset S_{\operatorname{{\mathop{\rm lcm}\nolimits}}(N^{\prime},M^{\prime})}$.
4.4. Pairing on $\sigma$-constructions
@@
For $N$, $N^{\prime}$ and $g$ as above, we have the commutative diagram
@@
Suppose $N$ is such that $N\geq 3$, $U_{N}\subset U$, the conductor of $\psi$ divides $N$. Then the morphism ${\mathcal{M}}_{F}\to{\mathcal{H}}$ induces an isomorphism
@@
We now construct the pairings by defining the duality homomorphisms. Given a representation $\sigma:U\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{{\mathcal{O}}_{K}}{\mathcal{V}}$, let $\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits})$ denote the representation defined by the action of $U$ on ${\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{V}},{\mathcal{O}}_{\psi})$. For $N\geq 3$ such that $U_{N}\subset\ker\sigma$ and the conductor of $\psi$ divides $N$, Lemma (4.7) and the duality homomorphism $\delta^{L}_{N}:{\mathcal{M}}_{N,c}\to{\mathop{\rm Hom}\nolimits}({\mathcal{M}}_{N},{\mathcal{T}}^{S_{N},\otimes(k-1)})$ from §2.5 give a morphism
@@
We will therefore omit the subscript $N$ in the notation for the morphisms $\bar{\delta}^{L}$. (Note however that $\delta^{L}_{N}$ depends on $N$, and $\delta^{L}_{N,!}$ depends on $N$ only in the normalization.)
@@
To do this, we give an alternate description of the Betti realizations and pairings which is independent of $N$. and define a perfect pairing on the new Betti realizations. Given a sufficiently small $U$ and a $\sigma:U\to\operatorname{{\mathop{\rm Aut}\nolimits}}_{{\mathcal{O}}_{K}}{\mathcal{V}}$, define a locally constant sheaf ${\mathcal{F}}(\sigma)$ on $G_{\mathbb{Q}}\backslash G_{\mathbb{A}}/UU_{\infty}$ by ${\mathcal{F}}(\sigma)=G_{\mathbb{Q}}\backslash(G_{\mathbb{A}}\times{\mathcal{V}})/UU_{\infty}$. Define
@@
4.5. Premotivic structure of level $N$ and character $\psi$
@@
for all ${\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in U_{0}(N)$. It is straightforward to check that this is precisely the space $M_{k}(N,\psi)$ of classical modular forms of weight $k$, level $N$ and character $\psi$, i.e., the set of forms $f\in M_{k}(\Gamma(M))$ such that
@@ 5.1. Hecke actions
Fix a rational prime $p$. By Lemma 4.2, $U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}_{p}U]_{\psi(p_{p})^{-1}}$ defines an endomorphism of ${\mathcal{M}}(N,\psi)_{M,!}$ regarded as an object in ${\mathcal{P}M}^{S_{pN}}$. This induces an endomorphism
@@ 5.2. Compatibility of the Hecke action with the pairings
Since $N$ is invertible in $\mathcal{O}$}, $UwU]_{\omega}$ is an isomorphism in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$ and
@@ 5.3. The action of Hecke rings
Fix a finite prime $p$. Recall that ${\mathcal{M}}(N,\psi)_{M,!}$ is in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$ and, by Lemma 4.2, $T_{p}:{\mathcal{M}}(N,\psi)_{M,!}\to{\mathcal{M}}(N,\psi)_{pM,!}$ is defined in ${\mathcal{P}M}^{S_{pM}^{K}}_{K}={\mathcal{P}M}^{S_{pN}^{K}}_{K}$. If $p|Nk!$, then $S_{pN}=S_{N}$. So $T_{p}$ is a morphism in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$.
@@ 6.3. $\Sigma$-variations for non-exceptional primes
In Proposition 4.4 take $U=U^{\Sigma_{0}},U^{\prime}=U^{\Sigma_{1}},U^{\prime\prime}=U^{\Sigma_{2}},g={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{m_{1}}}\end{array}\right)$}},g^{\prime}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{m_{2}}}\end{array}\right)$}}$ and $\tau=\tau^{\prime}={\mathop{\rm id}\nolimits}$. Then we have $U^{\prime}_{1}=U^{\prime}$. So in the double coset decomposition $U^{\prime}=\coprod_{k}U^{\prime}_{2}u_{k}U^{\prime}_{1}$ there is only one $u_{k}$ which we can take $I$. Further $W_{1}=W_{1}^{\prime}=U^{\prime\prime}$ and $\tau_{1}=1$. This gives us the desired equation.
@@
For this we first display the following elementary relations. Fix a prime number $p$. For each integer $i\geq 0$, let $V_{i}=\{{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})|c\equiv 0\mod p^{i}\}.$ Let $A_{t}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p^{t}}}\end{array}\right)$}},t=0,1,2$. Let $B_{0}=I,B_{1}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}},B_{2}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}$. Then
@@
Thus to prove Proposition 6.3, we only need to prove it in the case when $\Sigma$ contains one prime $p$. Let $\Sigma_{0}=\phi$ and $\Sigma_{1}=\Sigma$. We consider three cases depending on $\delta_{p}=0,1$ or $2$.
@@
When $q=p$, by (41) and (42), the double coset product of the local factor at primes other than $p$ involves $I$ and is obviously zero. The local factor at $p$ is
@@
We consider the action of $\gamma=\epsilon_{1}\circ\phi_{1}+\epsilon_{p}\circ\phi_{p}$ on $f$ as a classical modular form. First $\epsilon_{1}\circ\phi_{1}=[U{{}^{\Sigma}}IU]_{1}$ acts on ${\mathcal{M}}(N,\psi)_{M}\subseteq{\mathcal{M}}_{M}\otimes{\mathcal{V}}$ by $I\otimes 1$, where $M\geq 3$ is chosen such that $N|M$ and $S_{M}=S_{N}$. The action corresponding to $I$ on the classical forms is $I$. Since $f$ has rational coefficients, its image in $\oplus_{t\in({\mathbb{Z}}/M{\mathbb{Z}})^{\times}}M(M,\psi)$ is of the form $(f_{0}(z))_{t}$ with $f_{0}$ independent of $t$. Then, in the notations of Lemma 3.2 and Proposition 3.7, for each $t^{\prime}\in({\mathbb{Z}}/pM{\mathbb{Z}})^{\times}$ we have $\kappa(t^{\prime})=t^{\prime}\mod M$ and $\tilde{\gamma}:E_{pM,t^{\prime}}\to E_{M,t}$ defined by $\gamma=\gamma_{t^{\prime}}=I.$ So $(f_{0}(z))_{t}$ is sent to $(f_{0}(z))_{t^{\prime}}$, $t^{\prime}\in({\mathbb{Z}}/pM{\mathbb{Z}})^{\times}$ by $\epsilon_{1}\circ\phi_{1}$. Next $U{{}^{\Sigma}}=U{{}^{\Sigma}}\cap({\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}^{-1}$. So $\epsilon_{p}=p^{1-k}[U{{}^{\Sigma}}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}U]_{1}$ acts on ${\mathcal{M}}_{pM}\otimes{\mathcal{V}}$ by ${\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}\otimes 1$. The corresponding action on $\oplus_{t}M_{k}(\Gamma_{0}(pN))$ is given by taking, for each $t^{\prime}\in({\mathbb{Z}}/p^{2}M{\mathbb{Z}})^{\times}$, $\kappa(t^{\prime})=t^{\prime}\mod pM$ and $\tilde{\gamma}_{t^{\prime}}:E_{p^{2}M,t^{\prime}}\to E_{pM,t}$ given by $\gamma_{t^{\prime}}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p^{-1}}}\end{array}\right)$}}.$ So $\epsilon_{p}$ sends $(f_{0})_{t}\in\oplus_{t}M_{k}(\Gamma_{0}(pM))$ to $(f_{0}(pz))_{t^{\prime}}$. Since $\phi_{p}=-T_{p}$, $\epsilon_{p}\circ\phi_{p}$ sends $(f_{0}(z))_{t}$ to $(-f_{0}{}_{|T_{p}})_{t^{\prime}}$. So by [Miy, Lemma 4.6.5], $\gamma$ sends $(f_{0}(z))_{t}$ to $(f_{0}{{}^{\Sigma}}(z))_{t^{\prime}}$, as is needed.
@@
When $q\neq p$, the relation $T_{q}\gamma=\gamma T_{q}$ is proved in the same way as in Case 2. When $q=p$, we only need to consider the local factor of $T_{p}\gamma$ at the prime $p$. The double coset product of the local factor is
@@
As in the proof of Proposition 6.3, we can again specialize to the case when $\Sigma$ has a single prime $p$. Then we only need to consider the place at $p$.